# When isometry and equivalence for skew constacyclic codes coincide

### v1

@@ 4. Monomial homomorphisms between Petit algebras
We now turn to the question of when there exist monomial homomorphisms $G_{\tau,\alpha,k}$ from a Petit algebra $S[t;\sigma]/S[t;\sigma]f(t)$, where $f$ has degree $\mathbf{n}$, into the algebra $S[t;\sigma]/S[t;\sigma](t^{\mathbf{n}}-b)$.
It is not evident that $G$ is forced to be the reduction of a homomorphism $\tilde{G}:S[t;\sigma]\to S[t;\sigma]/S[t;\sigma](t^{\mathbf{n}}-b)$ as in Corollary 3.4. Namely, in $S[t;\sigma]/S[t;\sigma]f(t)$, the associativity of all powers of $t$ is a strong condition [Pet67, p. 13-06]: it is equivalent to $t^{\mathbf{n}}t=tt^{\mathbf{n}}$ and to $f(t)t\in S[t;\sigma]f(t)$. In particular, if all powers of $t$ are associative in $S[t;\sigma]/S[t;\sigma]f(t)$, then they are contained in the right nucleus of $S[t;\sigma]/S[t;\sigma]f(t)$, which is an associative subalgebra, and thus it is only in this case that Theorem 3.3 applies directly, as in the following example.
@@
Suppose $f(t)\in S_{0}[t]$. Then the associative algebra $S_{0}[t]/\langle f(t)\rangle$ is a subalgebra of $S[t;\sigma]/S[t;\sigma]f(t)$ containing all powers of $t$. It thus follows from Theorem 3.3 that if $G:S[t;\sigma]/S[t;\sigma]f(t)\to S[t;\sigma]/S[t;\sigma](t^{\mathbf{n}}-b)$ is a monomial homomorphism, then we must have
@@
When instead $S[t;\sigma]/S[t;\sigma]f(t)$ is proper nonassociative, then $t^{s}$ is only well-defined for $0\leq s<\mathbf{n}$, and the full power associativity of Theorem 3.3 is not required.
@@
$G_{\tau,\alpha}:S[t;\sigma]/S[t;\sigma](t^{\mathbf{n}}-a)\to S[t;\sigma]/S[t;\sigma](t^{\mathbf{n}}-b)$
@@
Suppose $f,g\in S[t;\sigma]$ are monic polynomials of degree $\mathbf{n}$. If for some cancellable $\alpha\in S$ and degree $k>1$, the map $G_{\tau,\alpha,k}:S[t;\sigma]/S[t;\sigma]f(t)\to S[t;\sigma]/S[t;\sigma]g(t)$ is a monomial homomorphism, then $k\equiv 1\mod\mathbf{m}$; if it is an isomorphism then $\gcd(k,\mathbf{n})=1$ and $\alpha\in S^{\times}$.
@@
$G=G_{\tau,\alpha,k}:S[t;\sigma]/S[t;\sigma](t^{\mathbf{n}}-a)\to S[t;\sigma]/S[t;\sigma](t^{\mathbf{n}}-b)$
@@
Let $\tau\in\operatorname{Aut}(S)$ commute with $\sigma$. Fix $\alpha\in S^{\times}$ and $2\leq k<\mathbf{n}$. Suppose first that the map $G=G_{\tau,\alpha,k}:S[t;\sigma]/S[t;\sigma](t^{\mathbf{n}}-a)\to S[t;\sigma]/S[t;\sigma](t^{\mathbf{n}}-b)$ defined by the rule
@@
Next, we note from Theorem 3.3 that for any $0<s<\mathbf{n}$, the element $G(t^{s})=G(t)^{s}\in S[t;\sigma]/S[t;\sigma](t^{\mathbf{n}}-b)$ will be a monomial of degree $sk]_{\mathbf{n}}$, the residue of $sk$ mod $\mathbf{n}$. In order for $G$ to further preserve the relation $t^{s}d=\sigma^{s}(d)t^{s}$ for all $d\in S$ we require $G(t^{s})\tau(d)=\tau(\sigma^{s}(d))G(t^{s})$, which will hold if and only if
@@
Since $\mathbf{m}|\mathbf{n}$ and $k\equiv 1\mod\mathbf{m}$, the relation $\alpha b=\sigma^{\mathbf{n}}(\alpha)\sigma^{k}(b)$ simplifies to $\sigma(b)=b$, or $b\in S_{0}$, yielding part of the condition (3).
@@
Finally, if $k,\mathbf{n},\mathbf{m},\alpha,a,b,\tau$ satisfy the first four conditions, then the $S_{0}^{\tau}$-linear map $G:S[t;\sigma]/S[t;\sigma](t^{\mathbf{n}}-a)\to S[t;\sigma]/S[t;\sigma](t^{\mathbf{n}}-b)$ given by (4.2) is a well-defined homomorphism of $S_{0}^{\tau}$-modules, and respects the relations $G(t^{\mathbf{n}})=G(a)$ and $G(t^{s}u)=G(\sigma^{s}(u)t^{s})$ for all $0<s<\mathbf{n}$ and all $u\in S$. Therefore in particular it satisfies $G(p)G(q)=G(pq)$ and $G(p+q)=G(p)+G(q)$ for all $p,q\in S[t;\sigma]/S[t;\sigma](t^{\mathbf{n}}-a)$, which implies it is an homomorphism of $S_{0}^{\tau}$-algebras. The final condition ensures it is an isomorphism. ∎
@@
Suppose $\mathbf{m}\nmid\mathbf{n}$, or that one of $a$ or $b$ is not in $S_{0}$. Then any Hamming-weight-preserving homomorphism
$G:S[t;\sigma]/S[t;\sigma](t^{\mathbf{n}}-a)\to S[t;\sigma]/S[t;\sigma](t^{\mathbf{n}}-b)$
@@
If $G$ is Hamming weight preserving, then it must send $t$ to an element of Hamming weight $1$, hence it is monomial. Further, the coefficient $\alpha$ in this case cannot be a zero divisor, or else there would be some $c\in S$ for which $G(ct)=0$. By Theorem 4.4, the degree of $G$ cannot be greater than one. ∎
When $a,b\in S_{0}$, on the other hand, Theorem 4.4 gives a plethora of monomial homomorphisms $G:S[t;\sigma]/S[t;\sigma](t^{\mathbf{n}}-a)\to S[t;\sigma]/S[t;\sigma](t^{\mathbf{n}}-b)$ of degree $k>1$.
@@
$S[t;\sigma]/S[t;\sigma](t^{\mathbf{n}}-r^{2}b^{3})\to S[t;\sigma]/S[t;\sigma](t^{\mathbf{n}}-b)$
@@ 5. Applications
In Section 4, we classified all monomial homomorphisms between Petit algebras of the form $S[t;\sigma]/S[t;\sigma](t^{\mathbf{n}}-a)$. Since the left ideals of these algebras give rise to skew constacyclic codes, we have classified the Hamming-weight preserving maps, or isometries, between their corresponding skew constacyclic codes.
@@ 5.1. Isometric constacyclic codes
Two rings $S[t;\sigma]/S[t;\sigma](t^{\mathbf{n}}-a)$ and $S[t;\sigma]/S[t;\sigma](t^{\mathbf{n}}-b)$ are called isometric if there exists a monomial isomorphism $G_{\tau,\alpha,k}$ between them for some $\alpha\in S^{\times}$ and $1\leq k<\mathbf{n}$. As a special case, we say they are equivalent if they are isomorphic via an isomorphism of the form $G_{\tau,\alpha}$ for some $\alpha\in S^{\times}$ and some $\tau\in\operatorname{Aut}(S)$ commuting with $\sigma$.
@@
Thus if two rings are $(\mathbf{n},\sigma)$-isometric, or $(\mathbf{n},\sigma)$-equivalent, then they are isometric, and if they are $(\mathbf{n},\sigma)$-equivalent then they are equivalent. Moreover, this definition captures all of the possible Hamming-weight preserving isomorphisms between these algebras, making it the correct notion of equivalence between codes.
@@ 5.2. A surprising consequence
Set $K=\mathbb{F}_{q}=\mathbb{F}_{p^{r}}$ with primitive element $\xi$, and let $\sigma(x)=x^{p^{s}}$. (Note that the order of $\sigma$ is $r/s$, not $s$, as stated.) Then $N_{\mathbf{n}}^{\sigma}(x)=x^{[\mathbf{n}]_{s}}$ for all $x\in K$ where
@@
In [OaNA25], $a,b\in S^{\times}$ are called $(\mathbf{n},\sigma)$-equivalent if $K[t;\sigma]/K[t;\sigma](t^{\mathbf{n}}-a)$ and $K[t;\sigma]/K[t;\sigma](t^{\mathbf{n}}-b)$ are $(\mathbf{n},\sigma)$-equivalent and $(\mathbf{n},\sigma)$-isometric if $K[t;\sigma]/K[t;\sigma](t^{\mathbf{n}}-a)$ and $K[t;\sigma]/K[t;\sigma](t^{\mathbf{n}}-b)$ are $(\mathbf{n},\sigma)$-isometric. Thus briefly, [OaNA25, Theorem 4] asserted the equivalence of:
@@
$K[t;\sigma]/K[t;\sigma](t^{\mathbf{n}}-a)$ and $K[t;\sigma]/K[t;\sigma](t^{\mathbf{n}}-b)$ are $(\mathbf{n},\sigma)$-isometric;
@@
$K[t;\sigma]/K[t;\sigma](t^{\mathbf{n}}-b^{k})$ and $K[t;\sigma]/K[t;\sigma](t^{\mathbf{n}}-a)$ are $(\mathbf{n},\sigma)$-equivalent, where $k$ is the degree of the isometry in (1). [^1] The actual conditions on $k$ specified were: $0<k<[\mathbf{m}]_{s}$, $\gcd(k,[\mathbf{m}]_{s})$.
@@
Choose the parameters $p,r,s,\mathbf{n}$ to satisfy that $(p^{r}-1)|[\mathbf{n}]_{s}$. For example, one may take $p=3$, $r=2$, $s=1$ and $\mathbf{n}=4$, so that $\mathbf{m}=2$ and $p^{r}-1=8$ (meaning $\xi^{8}=1$) and $\mathbf{n}]_{s}=40$. Then $\langle\xi^{[\mathbf{n}]_{s}}\rangle=\{1\}$ and in this case there are very few isomorphisms between these rings.
@@
Since $N_{\mathbf{n}}^{\sigma}(\alpha)=1$ for all $\alpha$, Proposition 4.2 implies that if there exists an monomial isomorphism of degree one $G_{\tau,\alpha}:K[t;\sigma]/K[t;\sigma](t^{\mathbf{n}}-a)\to K[t;\sigma]/K[t;\sigma](t^{\mathbf{n}}-b)$ we would have $\tau(a)=b$, which is false. Moreover, since $a\notin K_{0}$, condition Theorem 4.4[(3)] fails, so there are no monomial homomorphisms of degree $k$ (for any $k>1$). Thus $K[t;\sigma]/K[t;\sigma](t^{\mathbf{n}}-a)$ and $K[t;\sigma]/K[t;\sigma](t^{\mathbf{n}}-b)$ are not isometric, and (1) fails.
@@
We return to the setting of Example 5.2, where $b=a^{k}$. Then $\operatorname{id}:K[t;\sigma]/K[t;\sigma](t^{\mathbf{n}}-a^{k})\to K[t;\sigma]/K[t;\sigma](t^{\mathbf{n}}-b)$ is an isomorphism satisfying (4), but we have already shown that $K[t;\sigma]/K[t;\sigma](t^{\mathbf{n}}-a)$ and $K[t;\sigma]/K[t;\sigma](t^{\mathbf{n}}-b)$ are not isometric, so again (1) fails.

### v3

@@ 4. Monomial homomorphisms between Petit algebras
We now turn to the question of when there exist monomial homomorphisms $G_{\tau,\alpha,k}$ from a Petit algebra $\mathbb{S}_{f}$, where $f$ has degree $\mathbf{n}$, into the algebra $\mathbb{S}_{t^{\mathbf{n}}-b}$.
It is not evident that $G$ is forced to be the reduction of a homomorphism $\tilde{G}:S[t;\sigma]\to S[t;\sigma]/S[t;\sigma](t^{\mathbf{n}}-b)$ as in Corollary 3.4. Namely, in $\mathbb{S}_{f}$, the associativity of all powers of $t$ is a strong condition [Pet67, p. 13-06]: it is equivalent to $t^{\mathbf{n}}t=tt^{\mathbf{n}}$ and to $f(t)t\in S[t;\sigma]f(t)$. In particular, if all powers of $t$ are associative in $\mathbb{S}_{f}$, then they are contained in the right nucleus of $\mathbb{S}_{f}$, which is an associative subalgebra, and thus it is only in this case that Theorem 3.3 applies directly, as in the following example.
@@
Suppose $f(t)\in S_{0}[t]$. Then the associative algebra $S_{0}[t]/\langle f(t)\rangle$ is a subalgebra of $\mathbb{S}_{f}$ containing all powers of $t$. It thus follows from Theorem 3.3 that if $G:\mathbb{S}_{f}\to\mathbb{S}_{t^{\mathbf{n}}-b}$ is a monomial homomorphism, then we must have
@@
When instead $\mathbb{S}_{f}$ is proper nonassociative, then $t^{s}$ is only well-defined for $0\leq s<\mathbf{n}$, and the full power associativity of Theorem 3.3 is not required.
@@
$G_{\tau,\alpha}:\mathbb{S}_{t^{\mathbf{n}}-a}\to\mathbb{S}_{t^{\mathbf{n}}-b}$
@@
Suppose $f,h\in S[t;\sigma]$ are monic polynomials of degree $\mathbf{n}$. If for some cancellable $\alpha\in S$ and degree $k>1$, the map $G_{\tau,\alpha,k}:\mathbb{S}_{f}\to\mathbb{S}_{h}$ is a monomial homomorphism, then $k\equiv 1\mod\mathbf{m}$; if it is an isomorphism then $\gcd(k,\mathbf{n})=1$ and $\alpha\in S^{\times}$.
@@
$G=G_{\tau,\alpha,k}:\mathbb{S}_{t^{\mathbf{n}}-a}\to\mathbb{S}_{t^{\mathbf{n}}-b}$
@@
Let $\tau\in\operatorname{Aut}(S)$ commute with $\sigma$. Fix $\alpha\in S^{\times}$ and $2\leq k<\mathbf{n}$. Suppose first that the map $G=G_{\tau,\alpha,k}:\mathbb{S}_{t^{\mathbf{n}}-a}\to\mathbb{S}_{t^{\mathbf{n}}-b}$ defined by the rule
@@
Next, we note from Theorem 3.3 that for any $0<s<\mathbf{n}$, the element $G(t^{s})=G(t)^{s}\in\mathbb{S}_{t^{\mathbf{n}}-b}$ will be a monomial of degree $sk]_{\mathbf{n}}$, the residue of $sk$ mod $\mathbf{n}$. In order for $G$ to further preserve the relation $t^{s}d=\sigma^{s}(d)t^{s}$ for all $d\in S$ we require $G(t^{s})\tau(d)=\tau(\sigma^{s}(d))G(t^{s})$, which will hold if and only if
@@
Since $\mathbf{m}\mid\mathbf{n}$ and $k\equiv 1\mod\mathbf{m}$, the relation $\alpha b=\sigma^{\mathbf{n}}(\alpha)\sigma^{k}(b)$ simplifies to $\sigma(b)=b$, or $b\in S_{0}$, yielding part of the condition (3).
@@
Finally, if $k,\mathbf{n},\mathbf{m},\alpha,a,b,\tau$ satisfy the first four conditions, then the $S_{0}^{\tau}$-linear map $G:\mathbb{S}_{t^{\mathbf{n}}-a}\to\mathbb{S}_{t^{\mathbf{n}}-b}$ given by (4.2) is a well-defined homomorphism of $S_{0}^{\tau}$-modules, and respects the relations $G(t^{\mathbf{n}})=G(a)$ and $G(t^{s}u)=G(\sigma^{s}(u)t^{s})$ for all $0<s<\mathbf{n}$ and all $u\in S$. Therefore in particular it satisfies $G(p)G(q)=G(pq)$ and $G(p+q)=G(p)+G(q)$ for all $p,q\in\mathbb{S}_{t^{\mathbf{n}}-a}$, which implies it is an homomorphism of $S_{0}^{\tau}$-algebras. The final condition ensures it is an isomorphism. ∎
Note that when $\mathbf{m}\geq\mathbf{n}-1$, Theorem 4.4 shows the only isomorphisms between $\mathbb{S}_{t^{\mathbf{n}}-a}$ and $\mathbb{S}_{t^{\mathbf{n}}-b}$ are monomial of degree one, recovering the result in Theorem 2.3.
@@
Suppose $\mathbf{m}\nmid\mathbf{n}$, or that one of $a$ or $b$ is not in $S_{0}$. Then any Hamming-weight preserving homomorphism
$G:\mathbb{S}_{t^{\mathbf{n}}-a}\to\mathbb{S}_{t^{\mathbf{n}}-b}$
@@
If $G$ is Hamming-weight preserving, then it must send $t$ to an element of Hamming-weight $1$, hence it is monomial. Further, the coefficient $\alpha$ in this case cannot be a zero divisor, or else there would be some $c\in S$ for which $G(ct)=0$. By Theorem 4.4, the degree of $G$ cannot be greater than one. ∎
When $a,b\in S_{0}$, on the other hand, Theorem 4.4 gives a plethora of monomial homomorphisms $G:\mathbb{S}_{t^{\mathbf{n}}-a}\to\mathbb{S}_{t^{\mathbf{n}}-b}$ of degree $k>1$.
@@
$\mathbb{S}_{t^{4}-r^{2}b^{3}}\to\mathbb{S}_{t^{4}-b}$
@@ 5. Applications
In Section 4, we classified all monomial homomorphisms between Petit algebras of the form $\mathbb{S}_{t^{\mathbf{n}}-a}$. Since the left ideals of these algebras give rise to skew constacyclic codes, we have classified the Hamming-weight preserving maps, or isometries, between their corresponding skew constacyclic codes.
@@ 5.1. Isometric constacyclic codes
Two rings $\mathbb{S}_{t^{\mathbf{n}}-a}$ and $\mathbb{S}_{t^{\mathbf{n}}-b}$ are called isometric if there exists a monomial isomorphism $G_{\tau,\alpha,k}$ between them for some $\alpha\in S^{\times}$ and $1\leq k<\mathbf{n}$. As a special case, we say they are equivalent if they are isomorphic via an isomorphism of the form $G_{\tau,\alpha}$ for some $\alpha\in S^{\times}$ and some $\tau\in\operatorname{Aut}(S)$ commuting with $\sigma$.
Two classes ${\bf C}_{a}$ and ${\bf C}_{b}$ of skew $(\sigma,a)$-constacyclic codes of length $\mathbf{n}$, respectively skew $(\sigma,b)$-constacyclic codes of length $\mathbf{n}$, are called isometric, if there exists an isometry
$G_{\tau,\alpha,k}:\mathbb{S}_{t^{\mathbf{n}}-a}\to\mathbb{S}_{t^{\mathbf{n}}-b}$
and equivalent if there exists an isometry
$G_{\tau,\alpha}:\mathbb{S}_{t^{\mathbf{n}}-a}\to\mathbb{S}_{t^{\mathbf{n}}-b}.$
@@
Thus if two rings are $(\mathbf{n},\sigma)$-isometric, or $(\mathbf{n},\sigma)$-equivalent, then they are isometric, and if they are $(\mathbf{n},\sigma)$-equivalent then they are equivalent. Moreover, this definition captures all of the possible Hamming-weight preserving isomorphisms between these ambient rings, making it the correct notion of equivalence between codes, as the isometries $G_{\tau,\alpha,k}:\mathbb{S}_{t^{\mathbf{n}}-a}\to\mathbb{S}_{t^{\mathbf{n}}-b}$ preserve the dimension of the codes as well. This seems counterintuitive if $k>1$ so let us give a clarifying example.
Example 5.2.
In the setting of Example 4.6, choose $w,b\in S_{0}$ to satisfy $b^{3}=w^{4}$. Then $G=G_{\operatorname{id},1,3}:\mathbb{S}_{t^{4}-b^{3}}\to\mathbb{S}_{t^{4}-b}$ is a monomial isomorphism of degree $3$. In this case, it is easy to check that twt-w is a right divisor of $t^{4}-b^{3}=(t^{3}+wt^{2}+w^{2}t+w^{3})(t-w)$ in $S[t;\sigma]$, so it generates an ideal $C$ (that is, a skew $(4,\sigma)$-constacyclic code) of dimension 41=34-1=3.
On the other hand, $G(t-w)=t^{3}-w$ is not a divisor of $t^{4}-b$ in $S[t;\sigma]$ but (as one computes directly) $G(t^{3}+wt^{2}+w^{2}t+w^{3})G(t-w)=0$ in $\mathbb{S}_{t^{4}-b}$, as expected. Thus it would be erroneous to conclude that $\dim(G(C))=\mathbf{n}-\dim(G(t-w))$. Instead, we identify the polynomial $c(t)$ of least degree in $G(C)$ (by computing $t^{s}(t^{3}-w)$ for $s=0,1,2,\cdots$). In this case, we find that $G(C)$ is also generated by its skew-cyclic shift
$(-w^{-1}t)(t-w)=t-bw^{-1}\in\mathbb{S}_{t^{4}-b}=S[t;\sigma]/S[t;\sigma](t^{\mathbf{n}}-b).$
We verify this new choice of generator is indeed a right divisor of $t^{4}-b$ in $S[t;\sigma]$ since
$t^{4}-b=(t^{3}-bw^{-1}t^{2}-b^{2}w^{-2}t+b^{3}w^{-3})(t-bw^{-1}).$
Now suppose two classes ${\bf C}_{a}$ and ${\bf C}_{b}$ are isometric. Then $t^{n}-a$ and $t^{n^{\prime}}-b$ must have the same degree $\mathbf{n}$, so the skew constacyclic codes in ${\bf C}_{a}$ and ${\bf C}_{b}$ are $S$-submodules of $S^{\mathbf{n}}$. Every skew $(\sigma,a)$-constacyclic code $C$ in ${\bf C}_{a}$ is generated by a monic $g\in R$ that is a right divisor of $t^{\mathbf{n}}-a$ in $S[t;\sigma]$, therefore every skew $(\sigma,a)$-polycyclic code $C$ in ${\bf C}_{b}$ corresponds to a free (and cyclic) $S$-submodule $Rg/R(t^{\mathbf{n}}-a)$. We also know that $C$ has dimension $k=\mathbf{n}-\deg(g)$ (e.g., see [Pum25, Theorem 2.9 (i)]).
Lemma 5.3.
If two classes ${\bf C}_{a}$ and ${\bf C}_{b}$ are isometric via some monomial isomorphism $G_{\tau,\alpha,\ell}$ of degree $\ell$, then the skew $(\sigma,a)$-constacyclic codes in ${\bf C}_{a}$ have the same Hamming distance, dimension and length as the skew $(\sigma,a)$-constacyclic codes in ${\bf C}_{b}$.
Proof.
By assumption, there exists an isomorphism of nonassociative ambient Petit rings $G_{\tau,\alpha,\ell}:\mathbb{S}_{t^{\mathbf{n}}-a}\to\mathbb{S}_{t^{\mathbf{n}}-b}$. This immediately implies that the length of the codes in both classes is $\mathbf{n}$, and so the ring isomorphism $G_{\tau,\alpha,\ell}$ preserves the length of codes. Since the morphism is monomial, it sends monomials of $\mathbb{S}_{t^{\mathbf{n}}-a}$ to monomials of $\mathbb{S}_{t^{\mathbf{n}}-b}$ and so preserves Hamming distance.
Let $g$ be a right divisor of $t^{\mathbf{n}}-a$ in $S[t;\sigma]$; then the ideal generated by $g$ in $\mathbb{S}_{t^{\mathbf{n}}-a}$, call it $C$, is a skew $(\sigma,a)$-constacyclic code and all $(\sigma,a)$-constacyclic codes arise in this way [Pum25, Section 2.2]. The image of a right ideal of $\mathbb{S}_{t^{\mathbf{n}}-a}$ under the isomorphism $G_{\tau,\alpha,k}$ is a right ideal of $\mathbb{S}_{t^{\mathbf{n}}-b}$, since $G_{\tau,\alpha,\ell}(ug)=G_{\tau,\alpha,\ell}(u)G_{\tau,\alpha,\ell}(g)$ for all $u\in S[t;\sigma]/S[t;\sigma](t^{\mathbf{n}}-a)$. Thus we have a one-to-one correspondence of the codes in $\mathbf{C}_{a}$ and those in $\mathbf{C}_{b}$. Since the ring isomorphism $G_{\tau,\alpha,\ell}$ is in particular a $\tau$-semilinear bijective map between the free $S$-modules $S[t;\sigma]/S[t;\sigma](t^{\mathbf{n}}-a)$ and $S[t;\sigma]/S[t;\sigma](t^{\mathbf{n}}-b)$, it restricts to a $\tau$-semilinear bijective map between the free $S$-modules $S[t;\sigma]g/S[t;\sigma](t^{\mathbf{n}}-a)$ and $S[t;\sigma]G_{\tau,\alpha,\ell}(g)/S[t;\sigma](t^{\mathbf{n}}-b)$. Therefore both $C$ and $G_{\tau,\alpha,k}(C)$ must have the same dimension. ∎
Remark 5.4.
When $S$ is a finite ring we can also argue as follows to get some intuition on why the dimension of a skew constacyclic code will be by preserved also under isometries $G_{\tau,\alpha,\ell}$. Let $R=S[t;\sigma]$, for short. Every ring isomorphism $G_{\tau,\alpha,\ell}$ canonically yields a one-one correspondence between the left ideal $Rg/R(t^{\mathbf{n}}-a)$ and the left ideal $RG_{\tau,\alpha,\ell}(g)/R(t^{\mathbf{n}}-b)$, which means that $|Rg/R(t^{\mathbf{n}}-a)|=|RG_{\tau,\alpha,\ell}(g)/R(t^{\mathbf{n}}-b)|$ have the same number of elements. We know that both are free submodules of $S^{\mathbf{n}}$, that means $|Rg/R(t^{\mathbf{n}}-a)|=|S|^{k}$ and $|G_{\tau,\alpha,\ell}(g)/R(t^{\mathbf{n}}-b)|=|S|^{k^{\prime}}$ for some positive integers $k$ and $k^{\prime}$, so that $k=k^{\prime}$ and both must have the same dimension.
@@ 5.2. A surprising consequence
Set $K=\mathbb{F}_{q}=\mathbb{F}_{p^{r}}$ with primitive element $\xi$, and let $\sigma(x)=x^{p^{s}}$. (Note that the order of $\sigma$ is $\frac{r}{\gcd(r,s)}$, not $s$, as stated.) Then $N_{\mathbf{n}}^{\sigma}(x)=x^{[\mathbf{n}]_{s}}$ for all $x\in K$ where
@@
In [OaNA25], $a,b\in K^{\times}$ are called $(\mathbf{n},\sigma)$-equivalent if $\mathbb{K}_{t^{\mathbf{n}}-a}:=K[t;\sigma]/K[t;\sigma](t^{\mathbf{n}}-a)$ and $\mathbb{K}_{t^{\mathbf{n}}-b}:=K[t;\sigma]/K[t;\sigma](t^{\mathbf{n}}-b)$ are $(\mathbf{n},\sigma)$-equivalent and $(\mathbf{n},\sigma)$-isometric if $\mathbb{K}_{t^{\mathbf{n}}-a}$ and $\mathbb{K}_{t^{\mathbf{n}}-b}$ are $(\mathbf{n},\sigma)$-isometric. Thus briefly, [OaNA25, Theorem 4] asserted the equivalence of:
@@
$\mathbb{K}_{t^{\mathbf{n}}-a}$ and $\mathbb{K}_{t^{\mathbf{n}}-b}$ are $(\mathbf{n},\sigma)$-isometric;
@@
$\mathbb{K}_{t^{\mathbf{n}}-b^{k}}$ and $\mathbb{K}_{t^{\mathbf{n}}-a}$ are $(\mathbf{n},\sigma)$-equivalent, where $k$ is the degree of the isometry in (1). [^1] The actual conditions on $k$ specified were: $0<k<[\mathbf{m}]_{s}$, $\gcd(k,[\mathbf{m}]_{s})$.
@@
Choose the parameters $p,r,s,\mathbf{n}$ to satisfy that $(p^{r}-1)|[\mathbf{n}]_{s}$. For example, one may take $p=3$, $r=2$, $s=1$ and $\mathbf{n}=4$, so that $\mathbf{m}=2$ and $p^{r}-1=8$ (meaning $\xi^{8}=1$) and $\mathbf{n}]_{s}=40$. Then $\langle\xi^{[\mathbf{n}]_{s}}\rangle=\{1\}$ and in this case there are very few isomorphisms between these Petit rings.
@@
Since $N_{\mathbf{n}}^{\sigma}(\alpha)=1$ for all $\alpha$, Proposition 4.2 implies that if there exists an monomial isomorphism of degree one $G_{\tau,\alpha}:\mathbb{K}_{t^{\mathbf{n}}-a}\to\mathbb{K}_{t^{\mathbf{n}}-b}$ we would have $\tau(a)=b$, which is false. Moreover, since $a\notin K_{0}$, condition Theorem 4.4[(3)] fails, so there are no monomial homomorphisms of degree $k$ (for any $k>1$). Thus $\mathbb{K}_{t^{\mathbf{n}}-a}$ and $\mathbb{K}_{t^{\mathbf{n}}-b}$ are not isometric, and (1) fails.
@@
We return to the setting of Example 5.5, where $b=a^{k}$. Let $S=K$ be a field. Then $\operatorname{id}:\mathbb{K}_{t^{\mathbf{n}}-a^{k}}\to\mathbb{K}_{t^{\mathbf{n}}-b}$ is an isomorphism satisfying (4), but we have already shown that $\mathbb{K}_{t^{\mathbf{n}}-a}$ and $\mathbb{K}_{t^{\mathbf{n}}-b}$ are not isometric, so again (1) fails.